# From Logits to Hierarchies: Hierarchical Clustering made Simple

### v1

@@ Appendix A Code implementation of the L2H algorithm
In Figure 4, we provide a Python implementation of the L2H algorithm proposed in this work using standard scientific computing libraries (NumPy, SciPy). As stated in Section 3, our algorithm only requires the logits as input. It can be executed on the CPU even for large datasets, e.g., with a runtime of less than a minute for ImageNet-1K. Note that our procedure can be applied to any pre-trained unsupervised model to perform hierarchical clustering. Further, it can also be applied to logits from a supervised model to infer a hierarchy of classes. The output of the algorithm is a list, that contains the two groups that are merged at each step, hence characterizing the hierarchy. The aggregation function for computing the score per group is a design choice (as described in Section B.2) that can be viewed as a hyperparameter.
  {minted}[fontsize=]python import numpy as np from scipy.special import softmax
def L2H(logits): ””” L2H Algorithm. Args: logits: Logits from model (N x K) where N number of datapoints in the dataset and K is the number of clusters Returns: steps: Merging steps characterizing the hierarchy
””” # Number of clusters is equal to size of last dimension in the logits K = logits.shape[-1] # Initialize groups of clusters to single clusters groups = [(c,) for c in range(K)] # Initialize list of steps that characterize hierarchy steps = [] # Given the logits for the whole dataset, compute assignments and predicted probabilities softmaxed_logits = softmax(logits, axis=-1) assignments = np.argmax(softmaxed_logits, axis=-1) pred_probs = np.max(softmaxed_logits, axis=-1) for step in range(1, K): # Compute score for for each group (which chosen aggregation function) score_per_gr = for group in groups: score_per_gr[group] = sum([np.mean(pred_probs[assignments == c]) for c in group]) # Get the group with the lowest score (lsg), will be merged at this iteration lsg = min(score_per_gr, key=score_per_gr.get) # Get the logits for datapoints assigned to the lowest score group logits_lsg = logits[np.where(np.isin(assignments, lsg))[0]] # Reassign datapoints in lsg to only clusters not in lsg, # and re-compute predicted probabilities msm_logits_lsg = np.zeros_like(logits_lsg) cls_not_in_lsg = [c for c in range(K) if c not in lsg] cls_in_lsg = [c for c in range(K) if c in lsg] msm_logits_lsg[:, cls_not_in_lsg] = softmax(logits_lsg[:, cls_not_in_lsg], axis=-1) msm_logits_lsg[:, cls_in_lsg] = 0. reassignments = np.argmax(msm_logits_lsg, axis=-1) re_pred_probs = np.max(msm_logits_lsg, axis=-1) # Compute the total reassigned predicted probability per cluster and average across # clusters in each group.Then select the group with the highest average. re_pp_per_group = group: np.mean([np.sum(re_pred_probs[reassignments == c]) for c in group]) for group in groups if group != lsg mtg = max(re_pp_per_group, key=re_pp_per_group.get) # Merge ‘lsg‘ with ‘mtg‘ and update ‘groups‘. groups = [gr for gr in groups if gr not in [lsg, mtg]] + [lsg + mtg] # Add merging in current iteration to steps steps.append((lsg, mtg)) return steps
@@ B.1 Datasets
In this work, we run experiments on four challenging vision datasets, namely CIFAR-10 and CIFAR-100 (Lake et al., 2015), Food-101 (Bossard et al., 2014) and ImageNet-1K (Deng et al., 2009). CIFAR-10 and CIFAR-100 are well-established object classification datasets. The CIFAR-10 dataset consists of 60000 32x32 colored images divided in 10 classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, truck. The train/test splits contain 50000 and 10000 images respectively. Similarly, also the CIFAR-100 dataset consists of 60000 32x32 colored images. However, they are organized into 100 classes. In addition, the 100 classes are grouped into 20 superclasses. As for CIFAR-10, the train/test splits also contain 50000 and 10000 images respectively. The Food101 dataset is a fine-grained classification dataset of food images, consisting of 101000 images for 101 classes. Images are high-resolution, up to 512 pixels side length. Images are split between 75750 training samples and 25250 test images. The ImageNet-1K dataset, widely used in computer vision, consists of 1000 classes organized according to the WordNet hierarchy (Miller, 1995), with 1281167 training and 50000 test samples, respectively.
@@ B.2 Implementation details
For our hierarchical clustering experiments, to train the TURTLE and TEMI models on all considered datasets, we use the official code provided by the authors with recommended choices for hyperparameters (Gadetsky et al., 2024; Adaloglou et al., 2023). In particular, TEMI employs CLIPViTL/14 representations of the data, while TURTLE employs both CLIPViTL/14 and DINOv2 ViT-g/14 representations. For more details on TURTLE trained using two representation spaces, see the original paper (Gadetsky et al., 2024). We train both TEMI and TURTLE with a number of clusters $K$ equal to the true number of classes in each dataset. For each dataset, we train models on the training set, then report metrics on the test set. Note that the L2H algorithm takes as input logits from the training set to infer the hierarchy, while metrics that evaluate the quality of the hierarchy are computed on the test set. As the aggregation function $\Lambda$ in the L2H algorithm (see Section 3) to score each group we employ
$s(G)=\sum_{c\in G}\frac{1}{|\mathcal{D}^{c}|}\sum_{\bm{x}\in\mathcal{D}^{c}}g_{\theta}(\bm{x})$
which we find to work well in practice. However, other choices are possiblle. We implement TreeVAE (Manduchi et al., 2023) with their contrastive approach using the provided PyTorch codebase with corresponding defaults. The splitting criterion is set to the number of samples, an inductive bias that benefits this baseline method, since all datasets are balanced (Manduchi et al., 2023; Vandenhirtz et al., 2024). We set the number of clusters to $10$ for CIFAR-10 and to $20$ for the rest, due to the computational complexity, as seen in Table 2, as well as memory complexity, since every additional leaf adds a new decoder. DeepECT (Mautz et al., 2020) is also implemented using their provided codebase with the augmented version. Note that similar to the results shown in Manduchi et al. (2023), for colored datasets, DeepECT fails to grow trees, as they always collapse, indicating that DeepECT fails to find meaningful splits. We implement agglomerative clustering using the scikit-learn library (Pedregosa et al., 2011), and fit the model using PCA embeddings of the datasets with $50$ components and wards criterion (Ward, 1963) as the linkage method. Using the author’s original codebase, we further train Hyperbolic Hierarchical Clustering (Liu et al., 2019) on CLIP embeddings of the respective datasets. The authors do not describe how to retrieve cluster assignments using their method, so we follow the agglomerative clustering procedure and assume the leaves of the last $K$ tree nodes created to form a cluster, where $K$ corresponds to the chosen number of clusters.
@@ B.3 Metrics
Flat clustering To assess model performance in flat clustering, for each model we take the clustering at the level of the hierarchy where the number of clusters corresponds to the true number of classes $K$ in a given dataset. If the number of leaves at the leaf level of the hierarchy is smaller than $K$, as is the case for, e.g., TreeVAE and DeepECT on CIFAR-100, we consider the clustering at the leaf level. For flat clustering comparisons, we resort to well-established metrics, namely NMI, ARI, Accuracy, and Purity of the clusters (i.e., Leaf Purity). To compute accuracy and leaf purity, we resort to recent implementations in Gadetsky et al. (2024) and Manduchi et al. (2023), respectively.
Hierarchical clustering To assess the quality of a learned hierarchy, and compare the results of different models in hierarchical clustering, we resort to two metrics. Dendrogram Purity (DP), introduced in Kobren et al. (2017), extends the notion of leaf purity to evaluate the purity of hierarchical clusters, and was recently adopted to benchmark hierarchical clustering models (Manduchi et al., 2023). Following the notation of Kobren et al. (2017), let $\mathcal{C}^{*}$ denote the true $K$-clustering (i.e., true class labeling) of a dataset $\mathcal{D}$. Then define
$\mathcal{P}^{*}=\Bigl{\{}(x_{i},x_{j})\forall x_{i},x_{j}\in\mathcal{D},x_{j}\neq x_{j}\mid C^{*}(x_{i})=C^{*}(x_{j})\Bigr{\}}$
as the set of pairs of data points that belong to the same true cluster. Also let $C^{*}_{k}$ be the set of datapoints in cluster $k$. Dendrogram Purity (DP) is then defined for a hierchical clustering $\mathcal{H}$ as
@@
where $\text{LCA}(x_{1},x_{2})$ computes the least common ancestor node of datapoints $x_{1}$ and $x_{2}$ in $\mathcal{H}$, $\text{lvs}(z)$ returns the set of leaves of the sub-tree rooted at any internal node $z$ and $\text{pur}(S_{1},S_{2})=|S_{1}\cap S_{2}|/|S_{1}|$. One possible caveat of this metric is its high correlation with Leaf Purity: with a high leaf purity, most pairs of samples sharing the true label will inevitably fall into the same leaf. To address this, we introduce an additional metric for evaluation, namely Least Hierarchical Distance. With a similar notation as above we define
$\bar{\mathcal{P}}^{*}=\Bigl{\{}(x_{i},x_{j})\forall x_{i},x_{j}\in\mathcal{D},x_{j}\neq x_{j}\mid C^{*}(x_{i})=C^{*}(x_{j})\wedge l(x_{i};\mathcal{H})\neq l(x_{j};\mathcal{H})\Bigr{\}}$
@@
where $td(l_{1},l_{2})$ computes the number of edges in the shortest path that connects two leaves $l_{1},l_{2}$ in the tree defined by $\mathcal{H}$. Different from Dendrogram Purity, Least Hierarchical Distance only takes into consideration pairs of data points with the same true label that do not fall into the same leaf. Hence, it does not exhibit strong correlation with Leaf Purity, being more specific to the quality of the hiearchy, rather than influenced by the clustering at the leaf level.
Appendix C Additional visualizations

### v2

@@ Appendix A Code Implementation of the L2H Algorithm
In Figure 4, we provide a Python implementation of the L2H algorithm proposed in this work using standard scientific computing libraries (NumPy, SciPy). As stated in Section 3, our algorithm only requires the logits as input. It can be executed on the CPU even for large datasets, e.g., with a runtime of less than a minute for ImageNet. Note that our procedure can be applied to any pre-trained unsupervised model to perform hierarchical clustering. Further, it can also be applied to logits from a supervised model to infer a hierarchy of classes. We store the hierarchy as a list comprised of groups of clusters that are merged iteratively. The aggregation function for computing the score per group is a design choice (as described in Section B.2) that can be viewed as a hyperparameter.
import numpy as np
from scipy.special import softmax
[…]
L2H Algorithm.
Args:
logits: Logits from model (N x K) where N number of datapoints in the dataset
and K is the number of clusters
Returns:
steps: Merging steps characterizing the hierarchy
[…]
# Number of clusters is equal to size of last dimension in the logits
K = logits.shape[-1]
# Initialize groups of clusters to single clusters
groups = [(c,) for c in range(K)]
# Initialize list of steps that characterize hierarchy
steps = []
# Given the logits for the whole dataset, compute assignments and predicted probabilities
softmaxed_logits = softmax(logits, axis=-1)
assignments = np.argmax(softmaxed_logits, axis=-1)
pred_probs = np.max(softmaxed_logits, axis=-1)
for step in range(1, K):
# Compute score for for each group (which chosen aggregation function)
score_per_gr = {}
for group in groups:
score_per_gr[group] = sum([np.mean(pred_probs[assignments == c]) for c in group])
# Get the group with the lowest score (lsg), will be merged at this iteration
lsg = min(score_per_gr, key=score_per_gr.get)
# Get the logits for datapoints assigned to the lowest score group
logits_lsg = logits[np.where(np.isin(assignments, lsg))[0]]
# Reassign datapoints in lsg to only clusters not in lsg,
# and re-compute predicted probabilities
msm_logits_lsg = np.zeros_like(logits_lsg)
cls_not_in_lsg = [c for c in range(K) if c not in lsg]
cls_in_lsg = [c for c in range(K) if c in lsg]
msm_logits_lsg[:, cls_not_in_lsg] = softmax(logits_lsg[:, cls_not_in_lsg], axis=-1)
msm_logits_lsg[:, cls_in_lsg] = 0.
reassignments = np.argmax(msm_logits_lsg, axis=-1)
re_pred_probs = np.max(msm_logits_lsg, axis=-1)
# Compute the total reassigned predicted probability per cluster and average across
# clusters in each group.Then select the group with the highest average.
re_pp_per_group = {
group: np.mean([np.sum(re_pred_probs[reassignments == c]) for c in group]) for
group in groups if group != lsg
mtg = max(re_pp_per_group, key=re_pp_per_group.get)
# Merge lsg with mtg and update groups‘.
groups = [gr for gr in groups if gr not in [lsg, mtg]] + [lsg + mtg]
# Add merging in current iteration to steps
steps.append((lsg, mtg))
return steps
@@ B.1 Datasets
In this work, we run experiments on five challenging vision datasets, namely CIFAR-10 and CIFAR-100 (cifarpaper), Food-101 (Bossard:2014aa), ImageNet1K (imagenetpaper), and as well INaturalist21(van2021benchmarking) introduced in Appendix C. CIFAR-10 and CIFAR-100 are well-established object classification datasets. The CIFAR-10 dataset consists of 60000 32x32 colored images, divided in 10 classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, truck. The train/test splits contain 50000 and 10000 images respectively. Similarly, also the CIFAR-100 dataset consists of 60000 32x32 colored images. However, they are organized into 100 classes. In addition, the 100 classes are grouped into 20 superclasses. As for CIFAR-10, the train/test splits also contain 50000 and 10000 images respectively. The Food101 dataset is a fine-grained classification dataset of food images, consisting of 101000 images for 101 classes. Images are high-resolution, up to 512 pixels side length. Images are split between 75750 training samples and 25250 test images. The ImageNet-1k dataset, widely used in computer vision, consists of 1000 classes organized according to the WordNet hierarchy (wordnetpaper), with 1281167 training and 50000 test samples, respectively. The INaturalist21 dataset (van2021benchmarking) contains 2.7 million images of natural species labelled at different taxonomy levels.
@@ B.2 Implementation Details
For our hierarchical clustering experiments, to train the TURTLE and TEMI models on all considered datasets we use the official code provided by the authors with recommended choices for hyperparameters (gadetsky2024letlabelsunsupervisedtransfer; Adaloglou2023TEMI). In particular, TEMI employs CLIPViTL/14 representations of the data, while TURTLE employs both CLIPViTL/14 and DINOv2 ViT-g/14 representations. For more details on TURTLE trained using two representation spaces, see the original paper (gadetsky2024letlabelsunsupervisedtransfer). We train both TEMI and TURTLE with a number of clusters $K$ equal to the true number of classes in each dataset (unless explicitly stated otherwise, e.g. in the sensitivity analysis reported in Figure 5). For each dataset, we train models on the training set, then report metrics on the test set. Note that the L2H algorithm takes as input logits from the training set to infer the hierarchy, while metrics that evaluate the quality of the hierarchy are computed on the test set. As the aggregation function $\Lambda$ in the L2H algorithm (see Section 3) we employ
[…]
which we find to work well experimentally. However, other choices are possible (see also Table 7). We implement TreeVAE (manduchi2023tree) with their contrastive approach using the provided PyTorch codebase with corresponding defaults. The splitting criterion is set to the number of samples, an inductive bias that benefits this baseline method, since all datasets are balanced (manduchi2023tree; vandenhirtz2024sctree). We set the number of clusters to $10$ for CIFAR-10 and to $20$ for the rest, due to the computational complexity, as seen in Table 2, as well as memory complexity, since every additional leaf adds a new decoder. DeepECT (mautz2020deep) is also implemented using their provided codebase with the augmented version. Note that similar to the results shown in manduchi2023tree, for colored datasets, DeepECT fails to grow trees, as they always collapse, indicating that DeepECT fails to find meaningful splits. We implement agglomerative clustering using the scikit-learn library (scikit-learn), and fit the model using PCA embeddings of the datasets with $50$ components and Ward’s criterion (Ward1963HierarchicalGT) as the linkage method. Using the author’s original codebase, we further train Hyperbolic Hierarchical Clustering (liu2019hyperbolic) on CLIP embeddings of the respective datasets. The authors do not describe how to retrieve cluster assignments using their method, so we follow the agglomerative clustering procedure and assume the leaves of the last $K$ tree nodes created to form a cluster, where $K$ corresponds to the chosen number of clusters.
@@ B.3 Metrics
Flat clustering To assess model performance in flat clustering, for each model we take the clustering at the level of the hierarchy where the number of clusters corresponds to the true number of classes $K$ in a given dataset. If the number of leaves at the leaf level of the hierarchy is smaller than $K$, as is the case for, e.g., TreeVAE and DeepECT on CIFAR-100, we consider the clustering at the leaf level. For flat clustering comparisons, we resort to well-established metrics, namely NMI, ARI, Accuracy, and Purity of the clusters (i.e., Leaf Purity). To compute accuracy and leaf purity, we resort to recent implementations in (gadetsky2024letlabelsunsupervisedtransfer) and (manduchi2023tree), respectively.
Hierarchical clustering To assess the quality of a learned hierarchy, and compare the results of different models in hierarchical clustering, we resort to two metrics. Dendrogram Purity (DP), introduced in (dppaper), extends the notion of leaf purity to evaluate the purity of hierarchical clusters, and was recently adopted to benchmark hierarchical clustering models (manduchi2023tree). Following the notation of (dppaper), let $\mathcal{C}^{*}$ denote the true $K$-clustering (i.e., true class labeling) of a dataset $\mathcal{D}$. Then define
$\mathcal{P}^{*}=\Bigl\{(x_{i},x_{j})\forall x_{i},x_{j}\in\mathcal{D},x_{i}\neq x_{j}\mid C^{*}(x_{i})=C^{*}(x_{j})\Bigr\}$
as the set of pairs of data points that belong to the same true cluster. Dendrogram Purity (DP) is then defined for a hierchical clustering $\mathcal{H}$ as
@@
where $\text{LCA}(x_{1},x_{2})$ computes the least common ancestor node of data points $x_{1}$ and $x_{2}$ in $\mathcal{H}$, $\text{lvs}(z)$ returns the set of leaves of the sub-tree rooted at any internal node $z$, $\text{pur}(S_{1},S_{2})=|S_{1}\cap S_{2}|/|S_{1}|$, and $\mathcal{C}_{k}^{*}$ is the set of data points belonging to the true cluster $k$. One possible caveat of this metric is its high correlation with Leaf Purity: with a high leaf purity, most pairs of samples sharing the true label will inevitably fall into the same leaf. To address this, we introduce an additional metric for evaluation, namely Least Hierarchical Distance. With a similar notation as above we define
$\bar{\mathcal{P}}^{*}=\Bigl\{(x_{i},x_{j})\forall x_{i},x_{j}\in\mathcal{D},x_{i}\neq x_{j}\mid C^{*}(x_{i})=C^{*}(x_{j})\wedge l(x_{i};\mathcal{H})\neq l(x_{j};\mathcal{H})\Bigr\}$
@@
where $td(l_{1},l_{2})$ computes the number of edges in the shortest path that connects two leaves $l_{1},l_{2}$ in the tree defined by $\mathcal{H}$. Different from Dendrogram Purity, Least Hierarchical Distance only takes into consideration pairs of data points with the same true label that do not fall into the same leaf. Hence, it does not exhibit a strong correlation with Leaf Purity, being more specific to the quality of the hierarchy rather than influenced by the clustering at the leaf level.
Appendix C Additional Results and Visualizations
Table 3 presents the results of our L2H algorithm implemented on top of the TURTLE backbone model for hierarchical clustering on the ImageNet1K dataset (imagenetpaper). These results complement the ones shown in Table 1, demonstrating that our method achieves outstanding performance for hierarchical clustering, even on datasets with large scale and numerous classes. It is worth noting that a direct comparison with alternative approaches (e.g., DeepECT, TreeVAE) is not feasible in this setting, as these methods lack the scalability to handle datasets of this magnitude and complexity. The quantitative results underscore the effectiveness of our method, successfully uncovering a meaningful hierarchical structure in a dataset of this size and complexity.
NMI $(\uparrow)$ ARI $(\uparrow)$ ACC $(\uparrow)$ LP $(\uparrow)$ DP $(\uparrow)$ LHD $(\downarrow)$
L2H-TURTLE $0.882$ $0.621$ $0.726$ $0.744$ $0.560$ $0.210$
Table 3: Hierarchical clustering performance of our L2H method applied on top of the TURTLE pre-trained model on the ImageNet1K dataset.
In Table 4, we present results obtained by applying our proposed approach using an alternative backbone model to TEMI and TURTLE, namely the TCL flat clustering model (li2022tcl). Unlike TEMI and TURTLE, TCL does not rely on pre-trained foundation model representations. This ablation highlights the generality of our method, demonstrating that it can be effectively applied with diverse choices of the backbone model. Notably, our approach maintains good hierarchical clustering performance even when the underlying backbone exhibits weaker flat clustering capabilities than TEMI or TURTLE, still outperforming deep specialized hierarchical approaches (see Table 1).
Flat Hierarchical
NMI $(\uparrow)$ ARI $(\uparrow)$ ACC $(\uparrow)$ LP $(\uparrow)$ DP $(\uparrow)$ LHD $(\downarrow)$
CIFAR-10 L2H-TCL 0.785 0.744 0.868 0.877 0.733 0.398
CIFAR-100 L2H-TCL 0.547 0.215 0.343 0.437 0.218 0.351
Food-101 L2H-TCL 0.455 0.168 0.279 0.348 0117 0.396
Table 4: Results obtained applying our L2H method with the TCL model used as backbone on CIFAR-10, CIFAR-100, Food-101 datasets.
In this work, we apply our L2H algorithm on top of flat models (TURTLE and TEMI) that both leverage CLIP embeddings to cluster the data. Hence, for further comparisons in Table 5 we report the results obtained with agglomerative clustering with Ward’s linkage performed on pre-trained CLIP embeddings from CIFAR-10, CIFAR-100, and Food-101 datasets. While leveraging pre-trained CLIP embeddings improves the performance of agglomerative clustering, a direct comparison with the performance of L2H-TEMI (see Table 1), which is also solely based on CLIP embeddings, shows that our method still markedly outperforms this baseline.
Flat Hierarchical
NMI $(\uparrow)$ ARI $(\uparrow)$ ACC $(\uparrow)$ LP $(\uparrow)$ DP $(\uparrow)$ LHD $(\downarrow)$
CIFAR-10 Agg. (CLIP embeddings) 0.799 0.724 0.805 0.826 0.716 0.442
CIFAR-100 Agg. (CLIP embeddings) 0.690 0.386 0.531 0.637 0.341 0.343
Food-101 Agg. (CLIP embeddings) 0.868 0.730 0.837 0.872 0.703 0.299
Table 5: Quantitative results for flat and hierarchical clustering performance of agglomerative clustering performed on CLIP embeddings.
To provide further comparisons that complement the results in Section 4.1, Table 6 presents results on the CIFAR-100 dataset comparing our proposed approach alongside two additional methods. First, we evaluate TreeVAE (manduchi2023tree) trained on CLIP embeddings of CIFAR-100, demonstrating that our method achieves substantially better performance. We include this comparison as both TEMI and TURTLE have access to CLIP embeddings for training. Hence this comparison shows that, even in a setup where deep specialized hierarchical approaches share access to foundation model embeddings, our proposed approach achieves significantly better performance with much higher efficiency. In particular, training TreeVAE in this setup requires hours on a GPU, while training L2H-TURTLE requires under two minutes. Furthermore, we include a comparison with the recent method introduced by lowe2024empiricalstudyclusteringunseen, which combines UMAP dimensionality reduction with Ward’s agglomerative clustering for zero-shot clustering with pre-trained embeddings. Since this approach relies on agglomerative clustering, it presents the drawback of not allowing inference on a separate test set. Hence, we both train and evaluate the this method on the test set, which results in an unfair advantage. Despite this, our method still significantly outperforms this additional baseline.
NMI $(\uparrow)$ ARI $(\uparrow)$ ACC $(\uparrow)$ LP $(\uparrow)$ DP $(\uparrow)$ LHD $(\downarrow)$
TreeVAE+CLIP 0.665 0.285 0.255 0.255 0.181 0.285
Lowe et al (2024) 0.753 0.499 0.616 0.664 0.445 0.303
L2H-TEMI 0.778 0.565 0.682 0.701 0.502 0.298
L2H-TURTLE 0.917 0.831 0.896 0.897 0.803 0.235
Table 6: Performance comparison on the CIFAR-100 dataset of our L2H approach, using TURTLE and TEMI as backbone models, with the method from Lowe et al (2024) and TreeVAE trained on CLIP embeddings. Hierarchical and flat clustering metrics are reported.
In Table 7, we provide an ablation that reports the results of L2H-TURTLE on the hierarchical clustering experiments from Section 4.1 with different choices for the aggregation function $\Lambda$ in the L2H algorithm. The results indicate that tweaks in the aggregation function alter performance, though without abrupt changes in the metrics. These results also motivate our designated choice of aggregation function—corresponding to the last row—which works well experimentally.
[…]
DP $(\uparrow)$ LHD $(\downarrow)$ DP $(\uparrow)$ LHD $(\downarrow)$ DP $(\uparrow)$ LHD $(\downarrow)$
L2H-TURTLE $\sum_{c\in G}\sum_{\bm{x}\in\mathcal{D}^{c}}g_{\theta}(\bm{x})$ 0.988 0.258 0.801 0.244 0.758 0.294
$\frac{1}{|G|}\sum_{c\in G}\frac{1}{|\mathcal{D}^{c}|}\sum_{\bm{x}\in\mathcal{D}^{c}}g_{\theta}(\bm{x})$ 0.988 0.248 0.793 0.283 0.751 0.335
$\sum_{c\in G}\frac{1}{|\mathcal{D}^{c}|}\sum_{\bm{x}\in\mathcal{D}^{c}}g_{\theta}(\bm{x})$ 0.988 0.277 0.803 0.235 0.758 0.297
Table 7: Results for hierarchical clustering, in terms of Dendrogram Purity and Least Hierarchical Distance, implementing the L2H algorithm with different choices for the aggregation function $\Lambda$, on top of the TURTLE model.
In Figure 5 we provide an ablation to test the sensitivity of our approach to the value of the hyperparameter $K$, corresponding to the number of leaves in the hierarchy and the number of clusters set for the pre-trained flat model. In particular, we report the results for L2H-TURTLE on the CIFAR-100 dataset varying the value of $K$ across the range $\{85,90,95,100,105,110,115\}$. Note that this range is symmetric around the true number of classes/clusters equal to $100$. Therefore, we both explore the case of over- and under-estimating the true number of clusters at the leaf level of the hierarchy. We report the results for both flat (NMI, ARI, ACC, LP) and hierarchical (DP, LHD) metrics. Best performance across all metrics is achieved when $K$ is set to the true number of clusters, while the performance gracefully degrades when $K$ is set to be an over- or under-estimated value. This demonstrates the robustness and stability of our approach with respect to this hyperparameter, which is particularly important in practical settings where the exact true number of classes is not known a priori. Finally, we find the log-normalized TURTLE model loss to be indicative of the true value of $K$, with the minimum value achieved when $K$ equals the true number of clusters. In practical settings, one can consider using this metric to select the value of $K$ when a value/proxy for the true number of clusters is not available.
Next, we perform an experiment to test whether, with datasets that are inherently hierarchical, implementing our approach can yield an advantage over flat clustering with the backbone model. To do so, we consider the INaturalist21 dataset (van2021benchmarking), which contains 2.7 million images of natural species labelled at different taxonomy levels (1103 families, 273 orders, 51 classes, 13 phylums, 3 kingdoms). We train TURTLE to model clusters at the more fine-grained family taxonomy level ($K_{family}=1103$). Then we implement our L2H procedure on top, and use the produced hierarchy to make clustering predictions at coarser taxonomy levels. For instance, $K_{order}-1$ steps from the end of the procedure (see Algorithm 1), we have a $K_{order}$-clustering of the data points, and we test its performance on the test set (at the order taxonomy level). Then, we train instances of TURTLE at each coarse taxonomy level, i.e., $K\in\{273,51,13,3\}$, and report the corresponding test set performance for comparison. Notably, training TURTLE at the fine-grained level, and using our L2H approach to construct a hierarchy and make predictions at more coarse levels, achieves better clustering performance than training separate TURTLE models at each taxonomy level. Note that this performance improvement also comes with much lower compute time, as only a single instance of the TURTLE model is trained (at the finest-grained taxonomy level). We report the results in Table 8.
Model Family Order Class Phylum Kingdom
nmi ari acc lp nmi ari acc lp nmi ari acc lp nmi ari acc lp nmi ari acc lp
TURTLE 0.552 0.052 0.140 0.373 0.512 0.075 0.172 0.562 0.498 0.101 0.203 0.827 0.514 0.244 0.268 0.893 0.479 0.461 0.663 0.873
L2H-TURTLE 0.552 0.052 0.140 0.373 0.517 0.097 0.181 0.572 0.497 0.123 0.229 0.797 0.515 0.294 0.385 0.877 0.561 0.562 0.734 0.920
Table 8: Comparison on the INaturalist21 dataset between the clustering performance of our L2H approach at coarse levels of the hierarchy, and the results obtained by training an instance of the backbone model at each coarse level. Results are averaged across five independent runs, bolding the best results and results that are statistically indistinguishable from the best.
Figure 5: Sensitivity analysis for L2H-TURTLE on the CIFAR-100 dataset with respect to the $K$ hyperparameter, which corresponds to the number of leaves in the hierarchy and the number of clusters set for training the pre-trained flat model (TURTLE in this case). Note that the true number of clusters is equal to $100$. Results for both flat (NMI, ARI, ACC, LP) and hierarchical (DP, LHD) metrics are included, with standard deviations across five independent runs reported as shaded areas around the line indicating mean values. We also include the log-normalized TURTLE model loss—reported in the rightmost plot in the bottom row—that proves to be indicative for model selection with respect to the $K$ hyperparameter.
Finally, we provide an ablation where we test a slightly different merging strategy for our L2H algorithm. In particular, instead of selecting the lowest-scoring group $G^{*}\in\text{argmin}_{G\in\mathcal{G}}s(G)$ and merging it with the most related group $G^{\dagger}\in\text{argmax}_{G\in\mathcal{G}\setminus\{G^{*}\}}\frac{1}{|G|}\sum_{c\in G}rp(c)$ (see details in Section 3), we directly consider all possible pairs of groups for merging, and merge the pair $G^{*},G^{\dagger}$ that results in the highest value for $\frac{1}{|G^{\dagger}|}\sum_{c\in G^{\dagger}}rp(c)$, where $rp(c)=\sum_{\bm{x}\in\mathcal{D}^{G^{*}},h_{\theta}^{m}(\bm{x};G^{*})=c}g_{\theta}^{m}(\bm{x};G^{*})\;$ (as also defined in Equation 2). We report the results with this alternative merging strategy in Table 9, only reporting hierarchical clustering metrics, as the merging strategy does not impact the flat level clustering. Results show that considering all pairs of groups for merging at each step, which also results in a computational overhead, does not bring benefits in performance.
DP $(\uparrow)$ LHD $(\downarrow)$
L2H-TURTLE (alt) 0.797 0.246
L2H-TURTLE 0.803 0.235
Table 9: Performance comparison on the CIFAR-100 dataset of our L2H approach, using TURTLE as backbone model, comparing the merging strategy described in Algorithm 1 with an alternative merging strategy where at each step all pairs of groups of clusters are considered for merging.